# SelfCite: Self-Supervised Alignment for Context Attribution in Large Language Models

### v2

@@ 3.3 Evaluation
  • Fine-tuned models: LongCite-8B and 9B released by Zhang et al. (2024), trained on LongCite-45K, fine-tuned from Llama-3.1-8B (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024), respectively. Additionally, we consider a baseline of finetuning LongCite-8B using SimPO with the NLI rewards which resembles Huang et al. (2024), with the details in Appendix E.
@@ 3.4 Main Results
To better contextualize the gains of our proposed reward, we additionally implement a variant of SimPO using NLI-based citation precision/recall rewards from Huang et al. (2024) by using the same training pipeline and initialization as our SimPO, modifying only the reward function (see details in Appendix E). As shown in row of SimPO w/ NLI Rewards, this baseline improves LongCite-8B on 3 out of 5 datasets, but is still consistently outperformed by SelfCite. This result highlights that while NLI-based rewards are helpful, our SelfCite reward provides a more accurate signal for optimizing citation quality.
@@ 5 Related Work
Recent work has explored various approaches to teaching language models to generate citations, including fine-tuning with direct human feedback or annotations (Nakano et al., 2021; Menick et al., 2022; Slobodkin et al., 2024), annotation/feedback from external models (Huang et al., 2024), and prompting-based methods (Gao et al., 2022, 2023b) to explicitly incorporate relevant retrieved documents. Given the high cost of human annotation, Zhang et al. (2024) introduced CoF (“Coarse to Fine”), an automated multi-stage pipeline that simulates human annotation. This approach leverages proprietary LLMs for query generation, chunk-level retrieval, and sentence-level citation extraction, achieving high citation quality through supervised fine-tuning. However, it depends on larger proprietary models two proprietary APIs—GLM-4 for the LLM and Zhipu Embedding-v2 for retrieval [^2] https://open.bigmodel.cn/pricing— with carefully designed prompting, effectively distilling the capabilities of these proprietary models into much smaller models in 8B/9B. In contrast, our SelfCite aims at completely eliminating the reliance on annotations for citation, either from human or proprietary APIs. Instead, our method enables a small 8B model to assess citation quality itself using self-supervised reward signal from context ablation, effectively self-improving without external supervision. We additionally provide Table 9 to contrast the key differences between SelfCite and prior papers in Appendix G.
@@ 6 Conclusion and Limitations
In this work, we introduced SelfCite, a self-supervised framework that aligns LLMs to generate more accurate, fine-grained citations by directly leveraging their own probabilities for necessity and sufficiency rewards through context ablation. With best-of-N sampling and preference optimization, SelfCite significantly improves citation correctness on the LongBench-Cite benchmark without requiring human annotation, offering a promising self-improving direction towards verifiable and trustworthy LLMs.
SelfCite also has limitations: 1) While achieving good results with SimPO, integrating other preference optimization or reinforcement learning (RL) algorithms, e.g., PPO (Schulman et al., 2017), remains under explored. However, as prior study (Mudgal et al., 2024) shows BoN closely approximates the upper-bound scores of RL, we mainly validate our reward design using BoN, following the established practices (Gao et al., 2023a; Lightman et al., 2024), and further verify it with a training-based method, SimPO. 2) The SelfCite framework assumes the output probabilities of LLMs are available, which may not work for some proprietary LLMs. 3) While we focus on self-supervision in preference optimization for improving the quality of existing citations generated by LLMs, better unsupervised ways to kick-start LLMs’ ability in generating structured citations can be further explored.
@@ Appendix E Baseline: SimPO with NLI Rewards
To provide a stronger fine-tuned baseline, we implement a SimPO variant that adopts NLI-based citation rewards, following the design proposed by  Huang et al. (2024). For fair comparison, we keep our full SelfCite SimPO training pipeline—initializing from LongCite-8B and training on the LongCite-45k dataset—and modify only the reward function as a controlled experiment. This NLI-based reward combines two components:
  • Citation Recall Reward: This measures whether the full set of cited sentences entails the model-generated statement. It is equivalent to the Citation Recall Reward proposed by Huang et al. (2024).
  • Citation Precision Reward: This estimates whether each cited sentence is necessary by ablating one sentence at a time and testing whether the remaining span still entails the statement. If entailment fails after removing a sentence, it indicates that the sentence contributes uniquely to the justification. To reduce latency, we ablate all sentences when the citation contains 5 or fewer; otherwise, we randomly sample 5 for ablation. When there are N ablations, each ablation makes a reward of $\frac{1}{N}$, and finally all ablations sum up to 1.0. It resembles the Citation Precision Reward proposed by Huang et al. (2024).
We make both rewards positive and capped at 1.0, effectively constructing preference pairs for SimPO. We do not consider the Correctness Recall Reward from Huang et al. (2024), because the LongCite-45k training set does not contain ground-truth answers. All entailment scores are computed using the public NLI model google/t5_xxl_true_nli_mixture [^7] https://huggingface.co/google/t5_xxl_true_nli_mixture.
@@ Appendix F Zero-shot Evaluation on Chunk-level Citation Benchmark ALCE
We additionally include the zero-shot evaluation on the chunk-level citation benchmark ALCE (Gao et al., 2023b) and report the results in Table 8. We find that our baseline model, LongCite-8B, although under a zero-shot setting (it is trained on sentence-level citation but test on chunk-level citations), already outperforms the prompting-based approach from Gao et al. (2023b) by a substantial margin in both citation recall and precision. Incorporating NLI-based rewards from Huang et al. (2024) into our SimPO training yields further improvements. Most notably, our method—SimPO with SelfCite rewards—achieves the best performance among models trained on the same LongCite-45k dataset.
The last row of the table presents the best result reported by Huang et al. (2024), who fine-tuned their model using supervised data. However, this setting is not directly comparable to ours for several reasons:
@@ Appendix G Comparison with Prior Studies
We further provide a comparison table in Table 9 to contrast the key differences between SelfCite and other prior studies on producing citations from LLMs. Among all methods, SelfCite is the only approach that supports sentence-level citation generation in a single pass, leverages preference optimization, and scales to 128K-token contexts—all without requiring additional supervision. In contrast, prior work such as ALCE (Gao et al., 2023b) and  Huang et al. (2024) use chunk-level citations for shorter context ($\leq$8K) and require prompt-based or supervised NLI signals. ContextCite (Cohen-Wang et al., 2024), while being sentence-level, relies on a computationally expensive (at least 32 inference calls) process for random context ablation and trains a linear model for estimating the importance scores. This comparison underscores the practical advantages and technical contributions of SelfCite.
@@
Huang et al. (2024) ✗ (chunk-level) ✗ (8K) NLI + ground truth

### v3

@@ 3.3 Evaluation
  • Fine-tuned models: LongCite-8B and 9B released by Zhang et al. (2024), trained on LongCite-45K, fine-tuned from Llama-3.1-8B (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024), respectively. Additionally, we consider a baseline of finetuning LongCite-8B using SimPO with the NLI rewards which resembles Huang et al. (2024a), with the details in Appendix E.
@@ 3.4 Main Results
To better contextualize the gains of our proposed reward, we additionally implement a variant of SimPO using NLI-based citation precision/recall rewards from Huang et al. (2024a) by using the same training pipeline and initialization as our SimPO, modifying only the reward function (see details in Appendix E). As shown in row of SimPO w/ NLI Rewards, this baseline improves LongCite-8B on 3 out of 5 datasets, but is still consistently outperformed by SelfCite. This result highlights that while NLI-based rewards are helpful, our SelfCite reward provides a more accurate signal for optimizing citation quality.
@@ 5 Related Work
Recent work has explored various approaches to teaching language models to generate citations, including fine-tuning with direct human feedback or annotations (Nakano et al., 2021; Menick et al., 2022; Slobodkin et al., 2024), rewards from external NLI models (Huang et al., 2024a, b), and prompting-based methods (Gao et al., 2022, 2023b) to explicitly incorporate relevant retrieved documents. Given the high cost of human annotation, Zhang et al. (2024) introduced CoF (“Coarse to Fine”), an automated multi-stage pipeline that simulates human annotation. This approach leverages proprietary LLMs for query generation, chunk-level retrieval, and sentence-level citation extraction, achieving high citation quality through supervised fine-tuning. However, it depends on larger proprietary models two proprietary APIs—GLM-4 for the LLM and Zhipu Embedding-v2 for retrieval [^2] https://open.bigmodel.cn/pricing— with carefully designed prompting, effectively distilling the capabilities of these proprietary models into much smaller models in 8B/9B. In contrast, our SelfCite aims at completely eliminating the reliance on annotations for citation, either from human or proprietary APIs. Instead, our method enables a small 8B model to assess citation quality itself using self-supervised reward signal from context ablation, effectively self-improving without external supervision. We additionally provide Table 9 to contrast the key differences between SelfCite and prior papers in Appendix G.
@@ 6 Conclusion and Limitations
We present SelfCite, a self-supervised framework for aligning large language models (LLMs) to generate more accurate and fine-grained citations. By leveraging LLMs’ own output probabilities, SelfCite computes necessity and sufficiency rewards through context ablation, enabling preference optimization without relying on external annotations from human or proprietary APIs. Applying such rewards in best-of-N (BoN) sampling and SimPO fine-tuning can significantly improve the citation correctness on the LongBench-Cite benchmark, offering a promising self-improving direction towards verifiable and trustworthy LLMs.
SelfCite also has limitations: 1) While achieving strong results with SimPO, integrating other preference optimization or reinforcement learning (RL) algorithms, e.g., PPO (Schulman et al., 2017), remains under explored. However, prior work (Mudgal et al., 2024) shows that BoN closely approximates the performance upper bound of RL, and we follow established practice (Gao et al., 2023a; Lightman et al., 2024) to mainly validate our rewards through BoN, and further verify it with SimPO fine-tuning. 2) SelfCite assumes access to model output probabilities, which may not be feasible for closed-source models. 3) While our framework improves the quality of citations already generated by LLMs, discovering unsupervised methods to kick-start LLMs’ ability in generating structured citations from scratch remains an important direction for future research.
@@ Appendix E Baseline: SimPO with NLI Rewards
To provide a stronger fine-tuned baseline, we implement a SimPO variant that adopts NLI-based citation rewards, following the design proposed by  Huang et al. (2024a). For fair comparison, we keep our full SelfCite SimPO training pipeline—initializing from LongCite-8B and training on the LongCite-45k dataset—and modify only the reward function as a controlled experiment. This NLI-based reward combines two components:
  • Citation Recall Reward: This measures whether the full set of cited sentences entails the model-generated statement. It is equivalent to the Citation Recall Reward proposed by Huang et al. (2024a).
  • Citation Precision Reward: This estimates whether each cited sentence is necessary by ablating one sentence at a time and testing whether the remaining span still entails the statement. If entailment fails after removing a sentence, it indicates that the sentence contributes uniquely to the justification. To reduce latency, we ablate all sentences when the citation contains 5 or fewer; otherwise, we randomly sample 5 for ablation. When there are N ablations, each ablation makes a reward of $\frac{1}{N}$, and finally all ablations sum up to 1.0. It resembles the Citation Precision Reward proposed by Huang et al. (2024a).
We make both rewards positive and capped at 1.0, effectively constructing preference pairs for SimPO. We do not consider the Correctness Recall Reward from Huang et al. (2024a), because the LongCite-45k training set does not contain ground-truth answers. All entailment scores are computed using the public NLI model google/t5_xxl_true_nli_mixture [^7] https://huggingface.co/google/t5_xxl_true_nli_mixture.
@@ Appendix F Zero-shot Evaluation on Chunk-level Citation Benchmark ALCE
We additionally include the zero-shot evaluation on the chunk-level citation benchmark ALCE (Gao et al., 2023b) and report the results in Table 8. We find that our baseline model, LongCite-8B, although under a zero-shot setting (it is trained on sentence-level citation but test on chunk-level citations), already outperforms the prompting-based approach from Gao et al. (2023b) by a substantial margin in both citation recall and precision. Incorporating NLI-based rewards from Huang et al. (2024a) into our SimPO training yields further improvements. Most notably, our method—SimPO with SelfCite rewards—achieves the best performance among models trained on the same LongCite-45k dataset.
The last row of the table presents the best result reported by Huang et al. (2024a), who fine-tuned their model using supervised data. However, this setting is not directly comparable to ours for several reasons:
@@ Appendix G Comparison with Prior Studies
We further provide a comparison table in Table 9 to contrast the key differences between SelfCite and other prior studies on producing citations from LLMs. Among all methods, SelfCite is the only approach that supports sentence-level citation generation in a single pass, leverages preference optimization, and scales to 128K-token contexts—all without requiring additional supervision. In contrast, prior work such as ALCE (Gao et al., 2023b) and  Huang et al. (2024a) use chunk-level citations for shorter context ($\leq$8K) and require prompt-based or supervised NLI signals. ContextCite (Cohen-Wang et al., 2024), while being sentence-level, relies on a computationally expensive (at least 32 inference calls) process for random context ablation and trains a linear model for estimating the importance scores. This comparison underscores the practical advantages and technical contributions of SelfCite.
@@
Huang et al. (2024a) ✗ (chunk-level) ✗ (8K) NLI + ground truth